OpenStudy (anonymous):

Jemma surveyed the students at her school to find out if they like the beach and/or the mountains for vacation. The table below shows the results of the survey: Like Beach Do Not Like Beach Total Like Mountains 36 28 64 Do Not Like Mountains 56 38 94 Total 92 66 158 If a student does not like the mountains, what is the probability that student also does not like the beach? 39.1% 40.4% 59.5% 60.8%

OpenStudy (anonymous):

@hype.child @marymiller @markaskingalexandria1 @maha2458 @loveeaj22 @zackabdi @help?!?!?! @water_boy @goformit100 @OGthug @KayBell15 @loveeaj22 @andrearchat @absolutemaster @ab60093 @Q345 @Wanderlust @haley15122 @hunnygirllxox @HelpMePlease1999 @help?!?!?! @Keshaun47 @gabalishc @Grayson1228 @Gabylovesyou @Keshaun47 @Keshaun47 @waleedo212 @Luigi0210 @lxrena @lxrena @Lotusflower_16 @loveeaj22 @LunyMoony help ill medal da 1 person ta get da answer @king_kong_killer @Leong @hinarauf80

OpenStudy (anonymous):

bruh lol @imqwerty

imqwerty (imqwerty):

total number of students who don like mountain+don like beach =56+38+94=188 total number of students=92+66+158=316 probability percentage =(total no. of favorable events/total number of events) x 100 =188/316 x 100 =59.49 ~ 59.5%

OpenStudy (anonymous):

thanks soo much your the best! @imqwerty

OpenStudy (anonymous):

do you mind if i ask another question

imqwerty (imqwerty):

np.. ask :)

OpenStudy (anonymous):

want me to post again so i can give another medal to you?

imqwerty (imqwerty):

lol i don need medals :) they r of no use.. post here itself :D

OpenStudy (anonymous):

The table shows the number of students in a school who like dodgeball and/or ping pong: Like Dodgeball Do Not Like Dodgeball Total Like Ping Pong 26 18 44 Do Not Like Ping Pong 24 32 56 Total 50 50 100 What is the relative frequency, by row, of students who like both dodgeball and ping pong? 0.22 0.40 0.48 0.59

OpenStudy (anonymous):

im really not good at this subject im not trying to be annoying lol but i might have another question

imqwerty (imqwerty):

The total number of students who like ping pong is 44. 26 of the 44 also like dodge ball. therefore the required frequency is 26/44. =0.59

OpenStudy (anonymous):

got it thankss! XD

imqwerty (imqwerty):

welcome :)

OpenStudy (anonymous):

one more? i promise last one lol @imqwerty

imqwerty (imqwerty):

ok lol

OpenStudy (anonymous):

A group of students were surveyed to find out if they like babysitting and/or mowing lawns to earn spending money. The results of the survey are shown below: 32 students like babysitting 21 students like babysitting but do not like mowing lawns 16 students like mowing lawns 12 students do not like babysitting or mowing lawns Four students created the tables below to represent the data. BT represents babysitting and ML represents mowing lawns. Emily Like BT Do Not Like BT Total Like ML 21 12 32 Do Not Like ML 11 5 16 Total 32 17 48 Will Like BT Do Not Like BT Total Like ML 5 11 16 Do Not Like ML 12 21 33 Total 17 32 49 Sarah Like BT Do Not Like BT Total Like ML 12 21 32 Do Not Like ML 5 11 16 Total 17 32 48 Maria Like BT Do Not Like BT Total Like ML 11 5 16 Do Not Like ML 21 12 33 Total 32 17 49 Which student's table is correct? Emily Will Sarah Maria

OpenStudy (anonymous):

@imqwerty ?

imqwerty (imqwerty):

m confused in the last one u see the data is all similar ....emily and sarah have the same data but only 2 columns r interchanged nd same is the case with wil nd maria..

OpenStudy (anonymous):

its ok i got the answer already lol

OpenStudy (anonymous):

thanks tho